OpenStudy (anonymous):

which of the following is the correct factoring of x^3+27? a)(x+3)(x+3)(x+3) b)(x+3)(x-3)(x-3) c)(x+3)(x^2+3x+9) d)(x+3)(x^2-3x-9) e)(x-3)(x^2-3x+9)

myininaya (myininaya):

none of the above

myininaya (myininaya):

x^3+27=(x+3)(x^2-3x+9) and also there was one of question of yours it was somehting like this x^2+10x-70=(x+9)(x+b) what is a+b you never said anything about a so i assumed you meant x^2+10x-70=(x-7)(x+10) so 10-7=3

OpenStudy (anonymous):

it should be: (x+3)(x^2 -3x +9)